OpenStudy (anonymous):

Anyone supper-duper good at explaining calc (A/B) and want to help me with several questions?

OpenStudy (anonymous):

and i officially can't spell...

hero (hero):

Define super-duper good

OpenStudy (anonymous):

SUPER DUPER GOOD = HERO

OpenStudy (anonymous):

well i just need to be able to understand it, like do people usually understand you?

OpenStudy (pokemon23):

saifoo this is the link http://openstudy.com/#/updates/4edc24e1e4b01b4ec4928501

OpenStudy (anonymous):

okay...well i'm just gonna ask my first question. if m is the slope of the curve xy=2 and n is the slope of the curve \[x^2-y^2=3\] then what is the point of intersection of the two curves?

hero (hero):

Is that all? You can do that without calculus

OpenStudy (anonymous):

hey now, its late. but yes thats it

hero (hero):

You can use substitution to find the values I get y = 2sqrt{3}/sqrt{5}

hero (hero):

you can substitute that back into the original equation to find the x value

hero (hero):

I get x = 2 to make y look prettier, rationalize the denominator

OpenStudy (anonymous):

okay.. well its a multiple choice question and the choices are m= -n mn= -1 m= n mn= 1 mn= -2

OpenStudy (anonymous):

i should have said that in the first place, sorry

myininaya (myininaya):

if the slope of that one curve is m and the slope of that other curve is n if they are perpendicular then they will have intersection when mn=-1

OpenStudy (anonymous):

it doesn't say if they are perpendicular or not

myininaya (myininaya):

m=-y/x and n=x/y m*n=-1

myininaya (myininaya):

the slopes are negative reciprocals

OpenStudy (anonymous):

oh! okay

OpenStudy (anonymous):

okay here's another one. the line y= 3x + k is tangent to the curve \[y= x^3\] then k is equal to... a) 1 or -1 b) 0 c) 3 or -3 d) 4 or -4 e) 2 or -2

myininaya (myininaya):

how do you find a tangent line to y=x^3?

myininaya (myininaya):

if y=x^3, what is y'?

OpenStudy (anonymous):

3x^2

myininaya (myininaya):

and say we wanted to find the tangent line to y=x^3 at x=a with slope 3 y'|=3a^2=3 => a=pm 1 and i said slope 3 since we are saying to find the tangent line to y=x^3 that has form y=3x+k so we have y=3x+k if x=1, then y=(1)^3=1 1=3(1)+k solve for k if x=-1, then y=(-1)^3=-1 -1=3(1)+k solve for k

OpenStudy (saifoo.khan):

myininya? please check the questions list.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

next question... two tangents that can be drawn from the point (3,5) t o the parabola y=x^2 have what slopes? 1 and 5 0 and 4 2 and 10 2 and -1/2 2 and 4